OpenStudy (anonymous):

Let A , B and C be non-coplanar unit vectors , equally inclined to one another at an angle 'theta' .If: A x B + B x C =pA +qB +rC , Find scalars p , q and r in terms of 'theta' .

OpenStudy (turingtest):

I don't think I can solve this, but I am commenting so I get tagged on this question.

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

we're given : \[A\cdot A = B\cdot B = C\cdot C = 1\] and \[A\cdot B = B\cdot C = C\cdot A= \cos \theta \]

ganeshie8 (ganeshie8):

Next suppose the volume of parallelopiped formed by these 3 vectors as edges is \(V\), then we have : \[A\cdot B\times C = B\cdot C\times A = C\cdot A\times B = V\]

ganeshie8 (ganeshie8):

using the given equation form 3 equations based on Volume and solve the system you will get answer interms of V and keep in mind that the volume can be expressed in terms of \(\theta \)

OpenStudy (mathmate):

@ganeshie8 The expression is not symmetric, \(A\times B+B\times C \equiv(A-C)\times B = pA+qB+rC\) which makes q=0. I've come up with long expressions of each of A, B and C in terms of \(\cos(\theta)\), but was hoping to see some solutions in terms of, perhaps, vector spaces.

ganeshie8 (ganeshie8):

Ahh rest of it should be easy by interpreting the scalar triple product as volume of parallelopiped The volume equals 0 if any two vectors are equal in the triple product

ganeshie8 (ganeshie8):

\[A \times B + B \times C =pA+qB +rC\] we form 3 equations by dotting this both sides by A,B,C one by one : \[A\cdot [A \times B + B \times C] =A\cdot [pA+qB +rC]\implies V = p +\cos t(q+r)\] \[B\cdot [A \times B + B \times C] =B\cdot [pA+qB +rC]\implies 0 = q + \cos t(r+p)\] \[C\cdot [A \times B + B \times C] =C\cdot [pA+qB +rC]\implies V = r + \cos t(p+q)\] 3 equations and 3 unknowns - we can solve them :)

OpenStudy (mathmate):

@ganeshie8 It does look easy! Thanks.